EPS Junkyard Arts, Inc., had earnings of $441,000 for the year. The company had 38,000 shares...

EPS

Junkyard Arts, Inc., had earnings of $441,000 for the year. The company had 38,000 shares of common stock outstanding during the year and issued 3,400 shares of $100 par value preferred stock. The preferred stock has a dividend of $9 per share. There were no transactions in either common or preferred stock during the year.

Determine the basic earnings per share for Junkyard Arts for the year. Round answer to two decimal places.
$per share

Issuing Stock

Work Place Products Inc., a wholesaler of office products, was organized on July 1 of the current year, with an authorization of 125,000 shares of 3% preferred stock, $40 par and 500,000 shares of $10 par common stock. The following selected transactions were completed during the first year of operations:

Journalize the transactions.

July. 1. Issued 130,000 shares of common stock at par for cash.

July. 1

July. 1 Issued 450 shares of common stock at par to an attorney in payment of legal fees for organizing the corporation.

July. 1

Aug. 7. Issued 19,500 shares of common stock in exchange for land, buildings, and equipment with fair market prices of $35,000, $194,000, and $44,000, respectively.

For a compound transaction, if an amount box does not require an entry, leave it blank.

Aug. 7

Sept. 20. Issued 25,000 shares of preferred stock at $52 for cash.

For a compound transaction, if an amount box does not require an entry, leave it blank.

Sept. 20
Answer #1

Answer-1-

Preferred dividend = 3,400 * 9 = $ 30,600

EPS = (Net income - preferred dividend) / Number of shares outstanding

                = ( 441,000 -30,600 ) / 38,000

                = 410,400 /38,000

                = $ 10.80 per share

Answer-2-

Date accounts & explanation Debit credit
Jul 1 Cash (130,000*10) 1,300,000
Common stock 1,300,000
(To record issue common stock)
Jul 1 Organisation expenses (450*10) 4,500
Common stock 4,500
(To record issue common stock for organisation exp)
Aug 7 Land 35,000
Building 194,000
Equipment 44,000
Common stock (19,500*10) 195,000
Paid in capital in excess of par value- Common stock 78,000
(To record exchange common stock with Land, building and equipment)
Sep 20 Cash (25000*52) 1,300,000
Preferred stock (25000*40) 1,000,000
Paid in capital in excess of par value-preferred stock

300,

000

(To record issue of

preferred stock)

Kindly give me thumbs up if u like my answer....Thanks!!!!